# HIRE: A Hybrid Learned Index for Robust and Efficient Performance under Mixed Workloads

### v1

@@ Abstract.
copyright: ccjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: ACM SIGMOD/PODS International Conference on Management of Data; May 31–June 05, 2026; Bengaluru, Indiaisbn: 978-1-4503-XXXX-X/2018/06
@@ 1. Introduction
Indexes are essential components of modern DBMSs, and their efficiency impacts overall database performance for both retrieval and update operations (silberschatz2011database). The rapid advancement of machine learning has driven the emergence of learned indexes, which integrate predictive models with traditional indexing techniques (kraska2018case).
Learned indexes employ approximation models to learn the cumulative distribution function (CDF) of data, predicting data locations rather than traversing fixed algorithmic paths. Early studies demonstrated exceptional search performance, suggesting their potential to replace traditional index structures (kipf2020radixspline; stoian2021towards; hadian2021shift). However, maintaining updatability remains a significant challenge. Dynamic data updates can induce model drift, degrading prediction accuracy or necessitating computationally expensive retraining, thereby reducing the efficiency of learned indexes.
To address these challenges, updatable learned indexes, such as FIT-ting Tree (galakatos2019fiting), ALEX (ding2020alex), PGM (ferragina2020pgm), and LIPP (wu2021updatable), have been proposed, accompanied by extensive analyses (ferragina2020learned; liu2024learned; sun2023learned; wongkham2022updatable; liu2025good), optimizations (zhang2024revisiting; zhang2024making; choi2024can; chockchowwat2022tuning; wang2025new; zhong2022learned), and applications (chesetti2024evaluating; schmidt2024predicate; chatterjee2024limousine; liang2024swix; lan2024fully; yang2023flirt). However, while several studies suggest that updatable learned indexes are “almost ready” for separated update and query workloads (wongkham2022updatable), our experimental results reveal that learned indexes still exhibit high tail latency and inefficient range query performance under mixed workloads. As shown in Figure 1, a comparison of state-of-the-art learned indexes with B+-tree under a mixed workload highlights three key limitations: (1) inconsistent performance across different workloads; (2) suboptimal performance for range queries; and (3) unstable query/update latency. A detailed discussion of these limitations is provided in Section 2.
The primary reason for these limitations is that updatable learned index designs often prioritize optimal performance for point queries under specific conditions, at the cost of worst-case inefficiency and instability. For example, PGM (ferragina2020pgm) employs an index-level buffer to mitigate model drift during insertions, which renders range queries highly inefficient due to the need to search multiple trees. Similarly, LIPP (wu2021updatable) sacrifices data locality and balanced structure to enhance model accuracy, achieving low-cost point lookups and insertions but compromising range query performance.
@@ 2. Background and Motivation
Merge, Split, Retrain
@@
PGM (ferragina2020pgm) Mult. Balanced Tree
@@
LIPP (wu2021updatable) Imbalanced Tree Hierarchical Model
Random Scan
with Sequential Scan
@@
SWIX (liang2024swix) Two-level Queue
@@
Merge, Split, Retrain
@@
B+-tree (comer1979ubiquitous) Balanced Tree Hierarchical Search Sequential Scan Inplace Inplace
@@
RBTree (guibas1978dichromatic) Balanced Tree Hierarchical Search Random Scan Inplace Inplace
@@
Hybrid Sequential Scan
@@
In contrast to traditional indexes like B+-tree, learned indexes employ machine learning models to approximate the data distribution and predict key locations (kraska2018case). Specifically, a model $F$ learns the cumulative distribution function (CDF) of the indexed keys, enabling it to predict a search bound for a given key. Traditional indexes, such as B+-tree, use a parameter‐based search bound, such as the fanout $f$, to guide searches. In contrast, learned indexes rely on an error‐based search bound $\epsilon$, representing the potential deviation between the predicted and actual key location. This deviation requires a “last‐mile search” to locate the key precisely. Crucially, the model $F$ must be strictly monotonic (i.e., $\forall k_{m}<k_{n},\ F(k_{m})<F(k_{n})$). Without strict monotonicity, the last‐mile search range could span the entire table $\mathcal{T}$, rather than being bounded by $\epsilon$.
@@
  • ALEX (ding2020alex), a pioneering updatable learned index, employs linear interpolation to approximate the data distribution. This linearization enhances the generalization of the linear model and reduces the search space to a reasonable $\epsilon$. The interpolated space also accommodates insertions without significantly affecting model accuracy. However, linearization reduces data locality. When the data distribution is highly non‐linear and cannot be approximated by a single linear model within the error bound $\epsilon$, ALEX recursively partitions the data, using multiple linear models. This top‐down partitioning may result in an unbalanced index structure. While ALEX excels in point queries and updates in dynamic settings, its suboptimal data locality and potential for an unbalanced structure impair range query performance and introduce instability in operational latency.
  • PGM (ferragina2020pgm) uses piecewise linear models to segment data and constructs an internal node layer from these segments. However, its update mechanism adopts a semi-dynamic approach, buffering updates before periodically merging them into existing segments, which triggers retraining. This two-stage approach requires searching both the buffer and data segments during queries, impairing data locality. Consequently, PGM prioritizes read-heavy workloads, achieving superior average query and update performance in such scenarios. However, this results in higher query and update tail latency in write-intensive environments.
  • LIPP (wu2021updatable) aims for perfect precision by reducing $\epsilon$ to zero. However, it requires top‐down node splitting for non‐linear data distributions, which may result in a highly unbalanced structure. LIPP mitigates the impact of insertions on model accuracy by reserving extra space within nodes. While this enables excellent point query and insertion performance, it significantly compromises range query efficiency and introduces high tail latency.
@@ 3.2. Model-accelerated Internal Nodes
Each internal node stores a list of keys from its child nodes along with their corresponding pointers. Unlike leaf nodes, which also need to support efficient scans, internal nodes access behave more like point lookups. To optimize internal nodes, we integrate the search efficiency of learned indexes with a log-based update mechanism inspired by LSM-trees (DBLP:journals/acta/ONeilCGO96). For rapid lookups, each node employs a linear model over its keys. To improve model precision and update efficiency, the slope of the linear model is adjusted, and keys are remapped to accommodate insertions when child nodes split. Furthermore, by batching updates rather than executing them individually, the log structure serves as a buffer, avoiding the overhead of immediate retraining and delaying model invalidation after new insertions. To ensure accurate searches, the update log is accessed during each search and is designed to be lightweight and cache-efficient. As will be detailed in Section 4.2, our log-based update mechanism minimizes data movement while supporting fast lookups.
@@ 3.3. Non-Blocking Cost-Driven Recalibration
Updates in HIRE can trigger two types of recalibration. While low-cost structural adjustments, such as splits and merges, can be performed immediately, the expensive retraining of model-based the leaf node presents a significant challenge, as a naive approach would block operations and lead to high tail latency. To mitigate high tail latency and operation blocking, we propose a non-blocking, cost-model-driven recalibration mechanism. This mechanism uses background threads to execute costly retraining and structural adjustments without stalling foreground operations. Concurrently, a built-in cost model proactively determines the optimal moment to trigger these tasks by balancing performance gains against their overhead. This approach enables HIRE to adapt to changes in data distribution while maintaining stable, low tail latency. The detailed procedures are discussed in Section 4.3.
@@ 4. HIRE Operations
This section describes the operations of HIRE, including queries, updates, recalibration, and bulk loading.
@@ 4.3.2. Model-based Leaf Retraining
The balanced tree structure of HIRE facilitates index modifications, such as the replacement of specific subtrees. This inherent structural property is crucial for minimizing interference with concurrent operations accessing other regions of the index. To avoid blocking operations related to subtree reconstruction, we leverage modern multi-threaded hardware to design a non-blocking retraining mechanism based on Read-Copy-Update (RCU) (mckenney2001read). This approach allows nodes to be updated in a copy and atomically replaced without impeding queries. All operations of HIRE are executed under the Linux kernel RCU mechanism, using rcu_read_lock(), and rcu_read_unlock(), and rcu_dereference() to safely access pointers. Once the modifications are complete, a pointer to the old version is safely and atomically swapped to point to the new version using synchronize_rcu() and rcu_assign_pointer(). This approach ensures that concurrent read operations can traverse the index without acquiring locks and without observing inconsistent, intermediate states of the data structure. Consequently, HIRE can update leaf nodes and their corresponding paths to internal nodes with minimal disruption to ongoing operations, thereby maintaining high availability and throughput.
@@ 4.3.3. Legacy Leaf Node Transformation
Figure 6. CDFs of FACE, COVID, GENOME and PLANET Datasets
Figure 7. Throughput of B+-tree on Different Fanout
@@ 4.4. Bulk Loading
Our bulk-loading process employs a bottom-up strategy, controlled by a key hyperparameter: the error tolerance $\delta$. This parameter enables localized optimization during data partitioning, allowing for more efficient organization of the index structure. Given an input dataset of $N$ ordered keys, $K=\{k_{i},k_{2},\cdots,k_{N}\}$, the construction process scans this sorted sequence to build the index level-by-level, beginning at the leaf layer. For each parent node of the leaf layer, we first populate it with an initial set of child nodes. Specifically, we generate the first $\frac{f}{4}$ leaf nodes according to the data partitioning rules of leaf retraining described in Section 4.3. The partitioning keys of these leaf nodes are then used to fit an initial linear regression model, denoted as $\mathcal{F}$. For each subsequent leaf node that can be formed under this parent, we introduce a $\delta$-bounded optimization. Instead of using a fixed approach, we define a candidate window of $\delta$ keys around the original partitioning key. The optimal partitioning key is chosen from this window to minimize its deviation from the parent’s regression model $\mathcal{F}$. The deviation for a candidate key $k$ with an actual rank of $\mathcal{R}(k)$ is defined as a distance $\mathcal{D}(k)=\left|\mathcal{F}(k)-\mathcal{R}(k)\right|$. The key that minimizes this distance $\mathcal{F}(k)$ is selected as the final partitioning key and appended to the parent node. The model $\mathcal{F}$ is next updated in an online fashion using Recursive Least Squares (RLS) (hayes1996statistical).
@@ 4.5. Discussion
Multi-dimensional Keys. While HIRE is designed as an ordered index for single-attribute numeric keys, it can support multi-attribute keys using a standard lexicographical mapping, similar to that of the B+-tree (comer1979ubiquitous). Specifically, a composite key $\textbf{k}=(k_{1},\cdots,k_{m})$ is transformed into a single 1D key using an order-preserving function $P(\textbf{k})$. HIRE can then be built on these mapped keys. This approach is highly efficient for queries on the leading attribute. For example, a range predicate on the first column $k_{1}\in(L,U)$ can be converted into a query for a single, contiguous interval $P(L,min_{k_{2},\cdots,k_{m}}),P(U,max_{k_{2},\cdots,k_{m}})]$ on the mapped keys. HIRE then completes the search with a single range scan.
Figure 8. Throughput on Lookup Queries
Figure 9. Throughput on Range Queries
Figure 10. Tail Latency on Different Datasets
@@ 5.1. Experimental Setup
Datasets and Workloads. We conduct experiments on six real-world datasets from SOSD (kipf2019sosd) and GRE (wongkham2022updatable). All datasets consist of 200 million 64-bit unsigned integer keys, each paired with a random 64-bit unsigned integer value for experimental purposes. The following three datasets are selected from SOSD: OSM (uniformly sampled from OpenStreetMap locations), FACE (upsampled from Facebook user ID data), and AMZN (Amazon book sales popularity data). For GRE, we select the datasets COVID (uniformly sampled Twitter IDs labeled with COVID-19), GENOME (human chromosome gene location pairs), and PLANET (planet ID data derived from OpenStreetMap). The CDF of keys for the FACE, COVID, GENOME, and PLANET datasets is shown in Figure 6.
@@
Indexes for Evaluation. We compare our proposed HIRE with three general-purpose learned indexes previously discussed in Section 2: ALEX, [^1] https://github.com/microsoft/ALEX PGM, [^2] https://github.com/gvinciguerra/PGM-index and LIPP [^3] https://github.com/curtis-sun/TLI/tree/main/competitors/lipp. Note that we exclude SWIX (liang2024swix) from our experiments because it is specifically designed for streaming data workloads and does not support the full set of operations required by general-purpose indexes. We also employ B+-tree [^4] https://github.com/tlx/tlx/tree/master/tlx/container as a baseline representing traditional indexing structures. To ensure a fair comparison, SIMD acceleration is enabled for all baselines.
@@
We note that these parameters could be further optimized using automatic tuning tools like OpenTuner (DBLP:conf/IEEEpact/AnselKVRBOA14), DDPG (DBLP:journals/corr/LillicrapHPHETS15), or LITune (wang2025new). However, as automated parameter tuning is orthogonal to the core contributions of this work, we leave such exploration for future work.
@@
Figure 11. Throughput on Different Match Rates
Figure 12. Throughput and Index Size on Different Datasizes
Figure 13. Index Build Time on Different Datasizes

### v2

@@ Abstract.
Learned indexes, machine learning for databases management, query processing
copyright: ccjournal: PACMMODjournalyear: 2026journalvolume: 4journalnumber: 1 (SIGMOD)article: 43publicationmonth: 2doi: 10.1145/3786657ccs: Information systems Point lookupsccs: Information systems Unidimensional range searchccs: Theory of computation Data structures and algorithms for data management
@@ 1. Introduction
Indexes are essential components of modern DBMSs, and their efficiency impacts overall database performance for both retrieval and update operations (Silberschatz et al., 2011). The rapid advancement of machine learning has driven the emergence of learned indexes, which integrate predictive models with traditional indexing techniques (Kraska et al., 2018).
Learned indexes employ approximation models to learn the cumulative distribution function (CDF) of data, predicting data locations rather than traversing fixed algorithmic paths. Early studies demonstrated exceptional search performance, suggesting their potential to replace traditional index structures (Kipf et al., 2020; Stoian et al., 2021; Hadian and Heinis, 2021). However, maintaining updatability remains a significant challenge. Dynamic data updates can induce model drift, degrading prediction accuracy or necessitating computationally expensive retraining, thereby reducing the efficiency of learned indexes.
To address these challenges, updatable learned indexes, such as FIT-ting Tree (Galakatos et al., 2019), ALEX (Ding et al., 2020), PGM (Ferragina and Vinciguerra, 2020), and LIPP (Wu et al., 2021), have been proposed, accompanied by extensive analyses (Ferragina et al., 2020; Liu et al., 2024; Sun et al., 2023; Wongkham et al., 2022; Liu et al., 2025), optimizations (Zhang et al., 2024b, a; Choi et al., 2024; Chockchowwat, 2022; Wang et al., 2025; Zhong et al., 2022), and applications (Chesetti and Pandey, 2024; Schmidt et al., 2024; Chatterjee et al., 2024; Liang et al., 2024; Lan et al., 2024; Yang et al., 2023). However, while several studies suggest that updatable learned indexes are “almost ready” for separated update and query workloads (Wongkham et al., 2022), our experimental results reveal that learned indexes still exhibit high tail latency and inefficient range query performance under mixed workloads. As shown in Figure 1, a comparison of state-of-the-art learned indexes with B+-tree under a mixed workload highlights three key limitations: (1) inconsistent performance across different workloads; (2) suboptimal performance for range queries; and (3) unstable query/update latency. A detailed discussion of these limitations is provided in Section 2.
The primary reason for these limitations is that updatable learned index designs often prioritize optimal performance for point queries under specific conditions, at the cost of worst-case inefficiency and instability. For example, PGM (Ferragina and Vinciguerra, 2020) employs an index-level buffer to mitigate model drift during insertions, which renders range queries highly inefficient due to the need to search multiple trees. Similarly, LIPP (Wu et al., 2021) sacrifices data locality and balanced structure to enhance model accuracy, achieving low-cost point lookups and insertions but compromising range query performance.
@@ 2. Background and Motivation
In contrast to traditional indexes like B+-tree, learned indexes employ machine learning models to approximate the data distribution and predict key locations (Kraska et al., 2018). Specifically, a model $F$ learns the cumulative distribution function (CDF) of the indexed keys, enabling it to predict a search bound for a given key. Traditional indexes, such as B+-tree, use a parameter‐based search bound, such as the fanout $f$, to guide searches. In contrast, learned indexes rely on an error‐based search bound $\epsilon$, representing the potential deviation between the predicted and actual key location. This deviation requires a “last‐mile search” to locate the key precisely. Crucially, the model $F$ must be strictly monotonic (i.e., $\forall k_{m}<k_{n},\ F(k_{m})<F(k_{n})$). Without strict monotonicity, the last‐mile search range could span the entire table $\mathcal{T}$, rather than being bounded by $\epsilon$.
@@
Merge, Split,
Retrain
@@
PGM (Ferragina and Vinciguerra, 2020)
Mult. Balanced
Tree
@@
LIPP (Wu et al., 2021) Imbalanced Tree Hierarchical Model
Random Scan with
Sequential Scan
@@
SWIX (Liang et al., 2024) Two-level Queue
@@
Merge, Split,
Retrain
@@
B+-tree (Comer, 1979) Balanced Tree Hierarchical Search Sequential Scan Inplace Inplace
@@
RBTree (Guibas and Sedgewick, 1978) Balanced Tree Hierarchical Search Random Scan Inplace Inplace
@@
Hybrid Sequential
Scan
@@
  • ALEX (Ding et al., 2020), a pioneering updatable learned index, employs linear interpolation to approximate the data distribution. This linearization enhances the generalization of the linear model and reduces the search space to a reasonable $\epsilon$. The interpolated space also accommodates insertions without significantly affecting model accuracy. However, linearization reduces data locality. When the data distribution is highly non‐linear and cannot be approximated by a single linear model within the error bound $\epsilon$, ALEX recursively partitions the data, using multiple linear models. This top‐down partitioning may result in an unbalanced index structure. While ALEX excels in point queries and updates in dynamic settings, its suboptimal data locality and potential for an unbalanced structure impair range query performance and introduce instability in operational latency.
  • PGM (Ferragina and Vinciguerra, 2020) uses piecewise linear models to segment data and constructs an internal node layer from these segments. However, its update mechanism adopts a semi-dynamic approach, buffering updates before periodically merging them into existing segments, which triggers retraining. This two-stage approach requires searching both the buffer and data segments during queries, impairing data locality. Consequently, PGM prioritizes read-heavy workloads, achieving superior average query and update performance in such scenarios. However, this results in higher query and update tail latency in write-intensive environments.
  • LIPP (Wu et al., 2021) aims for perfect precision by reducing $\epsilon$ to zero. However, it requires top‐down node splitting for non‐linear data distributions, which may result in a highly unbalanced structure. LIPP mitigates the impact of insertions on model accuracy by reserving extra space within nodes. While this enables excellent point query and insertion performance, it significantly compromises range query efficiency and introduces high tail latency.
@@ 3.2. Model-accelerated Internal Nodes
Each internal node stores a list of keys from its child nodes along with their corresponding pointers. Unlike leaf nodes, which also need to support efficient scans, internal nodes access behave more like point lookups. To optimize internal nodes, we integrate the search efficiency of learned indexes with a log-based update mechanism inspired by LSM-trees (O’Neil et al., 1996). For rapid lookups, each node employs a linear model over its keys. To improve model precision and update efficiency, the slope of the linear model is adjusted, and keys are remapped to accommodate insertions when child nodes split. Furthermore, by batching updates rather than executing them individually, the log structure serves as a buffer, avoiding the overhead of immediate retraining and delaying model invalidation after new insertions. To ensure accurate searches, the update log is accessed during each search and is designed to be lightweight and cache-efficient. As will be detailed in Section 4.2, our log-based update mechanism minimizes data movement while supporting fast lookups.
Figure 2. Structure of HIRE
@@ 3.3. Non-Blocking Cost-Driven Recalibration
Updates in HIRE can trigger two types of recalibration. While low-cost structural adjustments, such as splits and merges, can be performed immediately, the expensive retraining of the model-based leaf node presents a significant challenge, as a naive approach would block operations and lead to high tail latency. To mitigate high tail latency and operation blocking, we propose a non-blocking, cost-model-driven recalibration mechanism. This mechanism uses background threads to execute costly retraining and structural adjustments without stalling foreground operations. Concurrently, a built-in cost model proactively determines the optimal moment to trigger these tasks by balancing performance gains against their overhead. This approach enables HIRE to adapt to changes in data distribution while maintaining stable, low tail latency. The detailed procedures are discussed in Section 4.3.
@@ 14     Retrain $inner_{l}$ and $inner_{l}$ models, and remap $\leftarrow$ start thread do;
Figure 4. Updates of HIRE
4.2.1. Insertion of New Data
4.2.2. Deletion of Existing Data
@@ 4.3.2. Model-based Leaf Retraining
The balanced tree structure of HIRE facilitates index modifications, such as the replacement of specific subtrees. This inherent structural property is crucial for minimizing interference with concurrent operations accessing other regions of the index. To avoid blocking operations related to subtree reconstruction, we leverage modern multi-threaded hardware to design a non-blocking retraining mechanism based on Read-Copy-Update (RCU) (McKenney et al., 2001). This approach allows nodes to be updated in a copy and atomically replaced without impeding queries. All operations of HIRE are executed under the Linux kernel RCU mechanism, using rcu_read_lock(), and rcu_read_unlock(), and rcu_dereference() to safely access pointers. Once the modifications are complete, a pointer to the old version is safely and atomically swapped to point to the new version using synchronize_rcu() and rcu_assign_pointer(). This approach ensures that concurrent read operations can traverse the index without acquiring locks and without observing inconsistent, intermediate states of the data structure. Consequently, HIRE can update leaf nodes and their corresponding paths to internal nodes with minimal disruption to ongoing operations, thereby maintaining high availability and throughput.
@@ 4.4. Bulk Loading
Our bulk-loading process employs a bottom-up strategy, controlled by a key hyperparameter: the error tolerance $\delta$. This parameter enables localized optimization during data partitioning, allowing for more efficient organization of the index structure. Given an input dataset of $N$ ordered keys, $K=\{k_{i},k_{2},\cdots,k_{N}\}$, the construction process scans this sorted sequence to build the index level-by-level, beginning at the leaf layer. For each parent node of the leaf layer, we first populate it with an initial set of child nodes. Specifically, we generate the first $\frac{f}{4}$ leaf nodes according to the data partitioning rules of leaf retraining described in Section 4.3. The partitioning keys of these leaf nodes are then used to fit an initial linear regression model, denoted as $\mathcal{F}$. For each subsequent leaf node that can be formed under this parent, we introduce a $\delta$-bounded optimization. Instead of using a fixed approach, we define a candidate window of $\delta$ keys around the original partitioning key. The optimal partitioning key is chosen from this window to minimize its deviation from the parent’s regression model $\mathcal{F}$. The deviation for a candidate key $k$ with an actual rank of $\mathcal{R}(k)$ is defined as a distance $\mathcal{D}(k)=\left|\mathcal{F}(k)-\mathcal{R}(k)\right|$. The key that minimizes this distance $\mathcal{F}(k)$ is selected as the final partitioning key and appended to the parent node. The model $\mathcal{F}$ is next updated in an online fashion using Recursive Least Squares (RLS) (Hayes, 1996).
@@ 4.5. Discussion
Multi-dimensional Keys. While HIRE is designed as an ordered index for single-attribute numeric keys, it can support multi-attribute keys using a standard lexicographical mapping, similar to that of the B+-tree (Comer, 1979). Specifically, a composite key $\textbf{k}=(k_{1},\cdots,k_{m})$ is transformed into a single 1D key using an order-preserving function $P(\textbf{k})$. HIRE can then be built on these mapped keys. This approach is highly efficient for queries on the leading attribute. For example, a range predicate on the first column $k_{1}\in(L,U)$ can be converted into a query for a single, contiguous interval $P(L,min_{k_{2},\cdots,k_{m}}),P(U,max_{k_{2},\cdots,k_{m}})]$ on the mapped keys. HIRE then completes the search with a single range scan.
Figure 6. CDFs of FACE, COVID, GENOME and PLANET Datasets
@@ 5.1. Experimental Setup
Datasets and Workloads. We conduct experiments on six real-world datasets from SOSD (Marcus et al., 2020) and GRE (Wongkham et al., 2022). All datasets consist of 200 million 64-bit unsigned integer keys, each paired with a random 64-bit unsigned integer value for experimental purposes. The following three datasets are selected from SOSD: OSM (uniformly sampled from OpenStreetMap locations), FACE (upsampled from Facebook user ID data), and AMZN (Amazon book sales popularity data). For GRE, we select the datasets COVID (uniformly sampled Twitter IDs labeled with COVID-19), GENOME (human chromosome gene location pairs), and PLANET (planet ID data derived from OpenStreetMap). The CDF of keys for the FACE, COVID, GENOME, and PLANET datasets is shown in Figure 6.
@@
Indexes for Evaluation. We compare our proposed HIRE with three general-purpose learned indexes previously discussed in Section 2: ALEX, [^1] https://github.com/microsoft/ALEX PGM, [^2] https://github.com/gvinciguerra/PGM-index and LIPP. [^3] https://github.com/curtis-sun/TLI/tree/main/competitors/lipp Note that we exclude SWIX (Liang et al., 2024) from our experiments because it is specifically designed for streaming data workloads and does not support the full set of operations required by general-purpose indexes. We also employ B+-tree [^4] https://github.com/tlx/tlx/tree/master/tlx/container as a baseline representing traditional indexing structures. To ensure a fair comparison, SIMD acceleration is enabled for all baselines.
Figure 7. Throughput of B+-tree on Different Fanout
@@
We note that these parameters could be further optimized using automatic tuning tools like OpenTuner (Ansel et al., 2014), DDPG (Lillicrap et al., 2016), or LITune (Wang et al., 2025). However, as automated parameter tuning is orthogonal to the core contributions of this work, we leave such exploration for future work.
@@
Figure 8. Throughput on Lookup Queries